# Near-Optimal Experiment Design in Linear non-Gaussian Cyclic Models

### v1

@@ 1 Introduction
Learning causal relationships among variables of interest in a complex system is a central goal in empirical sciences, forming the foundation for prediction, intervention, and explanation (pearl2009causality, ). These relationships are typically represented by causal graphs in which an edge from variable $X$ to variable $Y$ indicates that $X$ is a direct cause of $Y$. Much of the existing literature on causal structure learning assumes that the underlying causal graph is a Directed Acyclic Graph (DAG). However, many natural and engineered systems include feedback mechanisms that give rise to cycles in their causal representations. Such cyclic structures emerge in equilibrium models, low-frequency temporal sampling of dynamical systems, and various biological networks (bongers2021foundations, ).
@@
Interventions, i.e., actively perturbing the system and observing the resulting distributional changes can allow us to learn the graphs even with cycles. The experiment design problem studies how to best design interventions in order to maximize the information gained about the causal structure. Again, unlike the case of DAGs, where a body of work on experiment design exist, the work in the cyclic setting mokhtarian2023unified is few and far in between. This is in part due to the fact that, unlike in DAGs, where an intervention on a subset of the vertices orients all the edges incident to them, in cyclic directed graphs, performing experiments in some cases, neither leads to learning the presence of edges nor orienting them (mokhtarian2023unified, ).
@@
  • Experiments show that our adaptive strategy outperforms other heuristic methods and closely matches the feedback vertex set (FVS) lower bound (see more details in Section 8.2).
@@ 3 Preliminaries and Problem Formulation
Consider a Structural Causal Model (SCM) $\mathfrak{C}:=(\mathbf{S},P_{\mathbf{e}})$, where $\mathbf{S}$ consists of $n$ structural equations and $P_{\mathbf{e}}$ is the joint distribution over exogenous noise terms (peters2017elements, ). In a linear SCM (LSCM), the structural equations take the following form:
@@
In interventional structure learning, we perform $K$ distinct experiments to reduce uncertainty over the underlying causal graph. Each experiment involves a perfect intervention on a single variable. We denote the set of intervention targets by $I=\{i_{1},i_{2},\dots,i_{K}\}\subseteq[n]$, where $i_{j}\in[n]$ indicates that the $j$-th experiment intervenes on variable $x_{i_{j}}$. In the single-intervention setting we consider, each experiment replaces the structural assignment of the target variable with a new exogenous noise term (cf. Equation 2). As a result, data from the $j$-th experiment is drawn from the interventional distribution $P(\mathbf{x}|\mathrm{do}(x_{i_{j}}))$.
@@ 4.1 Strongly Connected Components
Two vertices $u$ and $v$ in a directed graph $G$ are said to be strongly connected if there are directed paths from $u$ to $v$ and from $v$ to $u$. This defines an equivalence relation on the vertex set, whose equivalence classes are called strongly connected components (SCCs). Each SCC is a maximal subgraph in which every pair of vertices is strongly connected, and no additional vertex from $G$ can be added without violating this property. The collection of SCCs forms a partition of the vertex set of $G$. See (cormen2022introduction, , Section 22.5)
@@ 4.2 Matching in Bipartite Graphs
Define a bipartite graph $G_{b}=(\{r_{1},\dots,r_{n}\},[n],E)$, where each node $r_{i}$ represents the $i$-th row of the matrix $I-W_{\mathrm{ICA}}$, and an edge $(r_{i},j)\in E$ exists if and only if $W_{\mathrm{ICA}}[i,j]\neq 0$. Each valid permutation $\pi$ defines a perfect matching $r_{i}\mapsto\pi(i)$, and vice versa. Thus, distribution-equivalent graphs correspond to row permutations consistent with matchings in $G_{b}$ (see an example in Appendix A.4).
@@ 5.1 Recovering an Edge of the True Matching in the Bipartite Graph
Proposition 5.1 has a natural interpretation in the bipartite graph of 4.2. Recall that each perfect matching in the bipartite graph $G_{b}=(\{r_{1},\dots,r_{n}\},[n],E)$, with edges $(r_{i},j)\in E$ iff $W_{\mathrm{ICA}}[i,j]\neq 0$, corresponds to a valid permutation matrix $P_{\pi}$ defining a distribution-equivalent graph.
@@ 6 Adaptive Experiment Design
Let $I=\{i_{1},i_{2},\dots,i_{K}\}\subseteq[n]$ be the set of selected variables for intervention. For each $i_{j}\in I$, we collect the interventional data from the distribution $\mathbf{x}^{(i_{j})}\sim P(\mathbf{x}|\mathrm{do}(x_{i_{j}}))$, and perform ICA to estimate the corresponding interventional matrix $I-W^{(i_{j})}$. By comparing this estimate with the observational estimate $I-W_{\mathrm{ICA}}$, we recover the row of $x_{i_{j}}$ and hence the corresponding matching edge. We then eliminate all graphs in $\mathcal{G}_{\text{eq}}$ that are incompatible with this edge.
Let $\Omega:=\mathcal{G}_{\text{eq}}$ denote the initial equivalence class. For a given intervention set $I\subseteq[n]$ and a true graph $\phi\in\Omega$, let $\Omega^{(I,\phi)}\subseteq\Omega$ denote the set of all graphs that are compatible with the result of interventions on variables in $I$. The objective is to find the set $I$ that minimizes the size of the reduced equivalence class $|\Omega^{(I,\phi)}|$. In the absence of a prior on the true graph $\phi\in\Omega$, we consider a uniform prior over $\Omega$. This leads to an expected-size criterion for selecting the next intervention.
@@ 6.1 Adaptive Stochastic Optimization
Following the Adaptive Stochastic Optimization framework (golovin2011adaptive, ), we define:
@@
  • For a given realization $\phi\in\Omega$ and policy $\pi$, let $I(\pi,\phi)=\bigl\{\,i_{1},i_{2},\dots,i_{K}\bigr\}$ be the (random) set of variables on which $\pi$ intervenes before terminating (subject to a budget of $K$ interventions).
@@
    $f(I,\phi):=|\Omega|-|\Omega^{(I,\phi)}|,$ (3)
    where $I\subseteq[n]$ denotes the set of intervened variables, and $\Omega^{(I,\phi)}\subseteq\Omega$ is the set of graphs that agree with $\phi$ on the outcomes of all interventions in $I$.
@@
$F(\pi)=\mathbb{E}_{\Phi\sim\mathrm{Unif}(\Omega)}\left[f\bigl(I(\pi,\Phi),\Phi\bigr)\right].$ (4)
@@ 6.2 Adaptive Monotonicity and Submodularity
We now review some relavent definitions from the adaptive submodularity framework (golovin2011adaptive, ), and show that our reward function satisfies these properties.
@@
Let $f(I,\phi)$ be the number of graphs eliminated after performing interventions $I$ under realization $\phi$, as defined in Equation (3). Then $f$ is adaptive monotone and adaptive submodular with respect to the uniform prior distribution over realizations.
Let $F(\pi)=\mathbb{E}_{\Phi\sim\mathrm{Unif}(\Omega)}\left[f\bigl(I(\pi,\Phi),\Phi\bigr)\right]$ denote the expected number of eliminated graphs under an adaptive policy $\pi$, as defined in Equation (4). Then the adaptive greedy policy achieves a (11/e)(1-1/e)-approximation to the optimal value of $F$.
@@ 7 Estimating the Reward Function by Sampling
The existence of a polynomial-time approximate uniform sampler is ensured by the FPRAS of Jerrum, Sinclair, and Vigoda (jerrum2004polynomial, ). While their method offers theoretical guarantees, we also consider faster heuristic alternatives in practice. For the sake of analysis, it suffices to assume access to such a sampler.
@@ 8 Experimental Results
We evaluate the performance of our adaptive experiment design method for identifying causal graphs under the linear non-Gaussian structural causal model assumption. Experiments are conducted on synthetic data generated from Erdős–Rényi random directed graphs, where each possible directed edge (excluding self-loops) is included independently with a fixed probability. To ensure identifiability, we enforce that the resulting matrix IWI-W is invertible. We set the intervention budget $K=n$, where $n$ is the number of nodes, and measure the average number of interventions required for full causal identification. A detailed description of the algorithm along with its pseudocode can be found in Appendix E.1.
8.1 Experimental Setup and Theoretical Baseline
To isolate the role of the intervention strategy, we assume access to an ideal ICA procedure that returns a row-permuted and scaled version of the true matrix IWI-W. The adaptive method uses a bipartite representation of the graph and samples perfect matchings using two modes: exact, where all matchings are enumerated, and sample, where a fast greedy heuristic is used (Details in Appendix E.2). Although the latter is non-uniform, it scales well and yields excellent empirical results.
@@ 8.1 Experimental Setup and Theoretical Baseline
Figure 2: Comparison of intervention strategies under two matching sampling regimes. Our method (Adaptive) consistently performs close to the feedback vertex set (FVS) lower bound.
As a theoretical baseline, we compute the feedback vertex set (FVS) of each true graph—defined as a minimal set of nodes whose removal makes the graph acyclic. Any experimenter, even with full knowledge of the graph, must intervene on at least one node in every cycle, so the FVS size represents a fundamental lower bound on the number of required interventions. Although computing the FVS is NP-hard, we solve it exactly via integer programming to obtain the best possible benchmark. Remarkably, our algorithm, without knowing the graph, operates in polynomial time (assuming ideal ICA) and performs consistently close to this intractable lower bound.
8.2 Comparison of Strategies
We compare our adaptive strategy against two baselines: Random, which selects an intervention target uniformly at random, and Max Degree, which chooses the node with the highest degree in the bipartite graph. Figures 2(a) and 2(b) show the average number of interventions under each strategy. The shaded region beneath the Adaptive curve represents the interval from $0.8\cdot\text{Adaptive}$ to Adaptive, illustrating how closely it tracks the FVS lower bound. Our method consistently requires fewer interventions than the baselines and often nearly matches the FVS, despite lacking access to the true graph structure. More simulations and details are presented in the appendix E.2.
@@ Appendix A Distribution Equivalence: Formal Definitions, Characterization, and Proofs
To understand how graphs within the same equivalence class differ, we introduce the cycle decomposition of permutations and relate it to structural changes in the graph. While this idea is discussed in [19], we include it here for completeness—especially since their treatment is limited to distribution-entailment equivalent LSCMs, rather than distribution-equivalent graphs.
@@ A.4 Example: Matchings in Bipartite Graphs
Here, cross marks indicate non-zero entries in the matrix, and the red entries mark a specific row permutation $\pi$ that selects a unique non-zero element in each row—corresponding to a perfect matching. This selection defines a valid permutation matrix $P_{\pi}$ such that $P_{\pi}(I-W_{\mathrm{ICA}})$ is a valid causal model in the equivalence class with non-zero elements on the diagonal. Figure 3(a) further illustrates how such permutations correspond to bipartite matchings.
@@ B.2 Example: How Interventions Reveal True Matching Edges
As discussed in Proposition 5.1, interventions provide critical information about the true matching in the bipartite representation. Recall that each perfect matching in the bipartite graph $G_{b}=(\{r_{1},\dots,r_{n}\},[n],E)$, where an edge $(r_{i},j)\in E$ exists iff $W_{\mathrm{ICA}}[i,j]\neq 0$, corresponds to a valid row permutation matrix $P_{\pi}$ defining a distribution-equivalent graph.
@@ Appendix C Proofs of Section 6
$f(I,\phi)\;=\;|\Omega|\;-\;\bigl|\Omega^{(I,\phi)}\bigr|$
@@
$I_{1}=\mathrm{dom}(\psi),\quad I_{2}=\mathrm{dom}(\psi^{\prime}),\quad I_{1}\subseteq I_{2},$
@@
$\Omega_{1}\;=\;\Omega^{(I_{1},\phi)},\quad\Omega_{2}\;=\;\Omega^{(I_{2},\phi)},$
@@
and $N_{2}\leq N_{1}$. Let $v\notin I_{2}$ be the next intervention candidate. In the bipartite‐matching view, assume that $v$ has $k$ possible edges across $\Omega_{1}$. Concretely, recall that each candidate graph in our current equivalence class $\Omega_{1}$ corresponds to a perfect matching in the bipartite graph $G_{b}$ whose left nodes are rows $r_{1},\dots,r_{n}$ and whose right nodes are column positions $1,\dots,n$. Before intervening on variable $v$, the right‐hand node $v$ may be matched to different rows $r_{i}$ across the $\lvert\Omega_{1}\rvert=N_{1}$ graphs. Call these distinct row‐indices $i=z_{1},z_{2},\dots,z_{k}$, so that the set of “candidate edges” for node $v$ is
@@
By the result of Golovin and Krause [11], adaptive monotonicity and submodularity guarantee that the adaptive greedy policy achieves at least a (11/e)(1-1/e)-approximation to the optimal policy. ∎

### v2

@@ 1 Introduction
Learning causal relationships among variables of interest in a complex system is a central goal in empirical sciences, forming the foundation for prediction, intervention, and explanation (pearl2009causality). These relationships are typically represented by causal graphs in which an edge from variable $X$ to variable $Y$ indicates that $X$ is a direct cause of $Y$. Much of the existing literature on causal structure learning assumes that the underlying causal graph is a Directed Acyclic Graph (DAG). However, many natural and engineered systems include feedback mechanisms that give rise to cycles in their causal representations. Such cyclic structures emerge in equilibrium models, low-frequency temporal sampling of dynamical systems, and various biological networks (bongers2021foundations).
@@
Interventions, i.e., actively perturbing the system and observing the resulting distributional changes can allow us to learn the graphs even with cycles. The experiment design problem studies how to best design interventions in order to maximize the information gained about the causal structure. Again, unlike the case of DAGs, where a body of work on experiment design exist, the work in the cyclic setting mokhtarian2023unified is few and far in between. This is in part due to the fact that, unlike in DAGs, where an intervention on a subset of the vertices orients all the edges incident to them, in cyclic directed graphs, performing experiments in some cases, neither leads to learning the presence of edges nor orienting them (mokhtarian2023unified).
@@
  • Experiments show that our adaptive strategy outperforms other heuristic methods and closely matches the feedback vertex set (FVS) lower bound (see more details in Section LABEL:sec:strategy-comparison).
@@ 3 Preliminaries and Problem Formulation
Consider a Structural Causal Model (SCM) $\mathfrak{C}:=(\mathbf{S},P_{\mathbf{e}})$, where $\mathbf{S}$ consists of $n$ structural equations and $P_{\mathbf{e}}$ is the joint distribution over exogenous noise terms (peters2017elements). In a linear SCM (LSCM), the structural equations take the following form:
@@
In interventional structure learning, we perform $K$ distinct experiments to reduce uncertainty over the underlying causal graph. Each experiment involves a perfect intervention on a single variable. We denote the set of intervention targets by $\mathcal{I}=\{i_{1},i_{2},\dots,i_{K}\}\subseteq[n]$, where $i_{j}\in[n]$ indicates that the $j$-th experiment intervenes on variable $x_{i_{j}}$. In the single-intervention setting we consider, each experiment replaces the structural assignment of the target variable with a new exogenous noise term (cf. Equation 2). As a result, data from the $j$-th experiment is drawn from the interventional distribution $P(\mathbf{x}|\mathrm{do}(x_{i_{j}}))$.
@@ 4.1 Strongly Connected Components
Two vertices $u$ and $v$ in a directed graph $G$ are said to be strongly connected if there are directed paths from $u$ to $v$ and from $v$ to $u$. This defines an equivalence relation on the vertex set, whose equivalence classes are called strongly connected components (SCCs). Each SCC is a maximal subgraph in which every pair of vertices is strongly connected, and no additional vertex from $G$ can be added without violating this property. The collection of SCCs forms a partition of the vertex set of $G$. See (cormen2022introduction, Section 22.5)
@@ 4.2 Matching in Bipartite Graphs
Define a bipartite graph $G_{b}=(\{r_{1},\dots,r_{n}\},[n],E)$, where each node $r_{i}$ represents the $i$-th row of the matrix $I-W_{\mathrm{ICA}}$, and an edge $(r_{i},j)\in E$ exists if and only if the $j$-th entry of row $r_{i}$ is non-zero. Each valid permutation $\pi$ defines a perfect matching $r_{i}\mapsto\pi(i)$, and vice versa. Thus, distribution-equivalent graphs correspond to row permutations consistent with matchings in $G_{b}$ (see an example in Appendix A.4).
@@ 5.1 Recovering an Edge of the True Matching in the Bipartite Graph
Proposition 5.1 has a natural interpretation in the bipartite graph of 4.2. Recall that each perfect matching in the bipartite graph $G_{b}=(\{r_{1},\dots,r_{n}\},[n],E)$, with edges $(r_{i},j)\in E$ iff the $j$-th entry of row $r_{i}$ is non-zero, corresponds to a valid permutation matrix $P_{\pi}$ defining a distribution-equivalent graph.
@@ 6 Adaptive Experiment Design
Let $\mathcal{I}=\{i_{1},i_{2},\dots,i_{K}\}\subseteq[n]$ be the set of selected variables for intervention. For each $i_{j}\in\mathcal{I}$, we collect the interventional data from the distribution $\mathbf{x}^{(i_{j})}\sim P(\mathbf{x}|\mathrm{do}(x_{i_{j}}))$, and perform ICA to estimate the corresponding interventional matrix $I-W^{(i_{j})}$. By comparing this estimate with the observational estimate $I-W_{\mathrm{ICA}}$, we recover the row of $x_{i_{j}}$ and hence the corresponding matching edge. We then eliminate all graphs in $\mathcal{G}_{\text{eq}}$ that are incompatible with this edge.
Let $\Omega:=\mathcal{G}_{\text{eq}}$ denote the initial equivalence class. For a given intervention set $\mathcal{I}\subseteq[n]$ and a true graph $\phi\in\Omega$, let $\Omega^{(\mathcal{I},\phi)}\subseteq\Omega$ denote the set of all graphs that are compatible with the result of interventions on variables in $\mathcal{I}$. The objective is to find the set $\mathcal{I}$ that minimizes the size of the reduced equivalence class $|\Omega^{(\mathcal{I},\phi)}|$. In the absence of a prior on the true graph $\phi\in\Omega$, we consider a uniform prior over $\Omega$. This leads to an expected-size criterion for selecting the next intervention.
@@ 6.1 Adaptive Stochastic Optimization
Following the Adaptive Stochastic Optimization framework (golovin2011adaptive), we define:
@@
  • For a given realization $\phi\in\Omega$ and policy $\pi$, let $\mathcal{I}(\pi,\phi)=\bigl\{\,i_{1},i_{2},\dots,i_{K}\bigr\}$ be the (random) set of variables on which $\pi$ intervenes before terminating (subject to a budget of $K$ interventions).
@@
    $f(\mathcal{I},\phi):=|\Omega|-|\Omega^{(\mathcal{I},\phi)}|,$ (3)
    where $\mathcal{I}\subseteq[n]$ denotes the set of intervened variables, and $\Omega^{(\mathcal{I},\phi)}\subseteq\Omega$ is the set of graphs that agree with $\phi$ on the outcomes of all interventions in $\mathcal{I}$.
@@
$F(\pi)=\mathbb{E}_{\Phi\sim\mathrm{Unif}(\Omega)}\left[f\bigl(\mathcal{I}(\pi,\Phi),\Phi\bigr)\right].$ (4)
@@ 6.2 Adaptive Monotonicity and Submodularity
We now review some relavent definitions from the adaptive submodularity framework (golovin2011adaptive), and show that our reward function satisfies these properties.
@@
Let $f(\mathcal{I},\phi)$ be the number of graphs eliminated after performing interventions $\mathcal{I}$ under realization $\phi$, as defined in Equation (3). Then $f$ is adaptive monotone and adaptive submodular with respect to the uniform prior distribution over realizations.
Let $F(\pi)=\mathbb{E}_{\Phi\sim\mathrm{Unif}(\Omega)}\left[f\bigl(\mathcal{I}(\pi,\Phi),\Phi\bigr)\right]$ denote the expected number of eliminated graphs under an adaptive policy $\pi$, as defined in Equation (4). Then the adaptive greedy policy achieves a (11/e)(1-1/e)-approximation to the optimal value of $F$.
@@ 7 Estimating the Reward Function by Sampling
The existence of a polynomial-time approximate uniform sampler is ensured by the FPRAS of Jerrum, Sinclair, and Vigoda (jerrum2004polynomial). While their method offers theoretical guarantees, we also consider faster heuristic alternatives in practice. For the sake of analysis, it suffices to assume access to such a sampler.
@@ 8 Experimental Results
We evaluate the performance of our adaptive experiment design method for identifying causal graphs under the linear non-Gaussian structural causal model assumption. Experiments are conducted on synthetic data generated from Erdős–Rényi random directed graphs, where each possible directed edge (excluding self-loops) is included independently with a fixed probability. To ensure identifiability, we enforce that the resulting matrix IWI-W is invertible. We measure the average number of interventions required for full causal identification.
8.1 Performance with an Ideal ICA Oracle
To evaluate the intervention strategy on its own, isolating it from statistical estimation errors, we assume access to an ideal ICA procedure that returns a row-permuted and scaled version of the true matrix IWI-W. The adaptive method uses a bipartite representation of the graph and samples perfect matchings using two modes: exact, where all matchings are enumerated, and sample, where a fast greedy heuristic is used (E.2). We compare our adaptive strategy against two baselines: Random, which selects a target uniformly at random, and Max Degree, which chooses the node with the highest degree in the bipartite graph.
As a theoretical baseline, we compute the feedback vertex set (FVS) of each true graph. The FVS size represents a fundamental lower bound on the number of interventions required to break all cycles. Although computing the FVS is NP-hard, we solve it exactly to obtain the best possible benchmark.
Figures 2(a) and 2(b) show that our adaptive method consistently outperforms the baselines and operates remarkably close to the intractable FVS lower bound, despite having no knowledge of the true graph structure. This demonstrates the near-optimality of our greedy approach in an ideal setting. For more details, please refer to E.
@@ 8.1 Performance with an Ideal ICA Oracle
Figure 2: Comparison of intervention strategies assuming an ideal ICA oracle. Our method (Adaptive) consistently performs close to the feedback vertex set (FVS) lower bound.
8.2 Robustness in a Practical Setting with Finite-Sample ICA
To assess the practical viability of our method, we evaluate its performance in a more realistic setting without an ideal ICA oracle. Instead, we apply the FastICA algorithm hyvarinen2000independent to finite samples generated from both observational and interventional distributions. This introduces estimation noise, which can corrupt the recovered matrices. To handle these challenges, we introduce two key algorithmic modifications for robustness: adaptive thresholding of matrix entries and a safe matching procedure to prevent incorrect row assignments. A detailed description of these modifications is provided in Appendix E.3.
In this setting, our adaptive strategy continues to exhibit a performance trend similar to the ideal case, consistently outperforming the Random and Max Degree baselines and remaining remarkably close to the FVS lower bound (see Figure 6 in Appendix E.3), confirming its effectiveness even under estimation noise of ICA.
Furthermore, to analyze the quality of the recovered structure, we measured the relative error ($\varepsilon_{\mathrm{rel}}=\frac{\left\|\widehat{W}-W^{*}\right\|_{F}}{\left\|W^{*}\right\|_{F}}$) between the estimated matrix $\widehat{W}$ and the ground-truth $W^{*}$. Our result, presented in Appendix E.3, shows that the algorithm achieves a high-fidelity recovery of the true causal structure in the vast majority of runs.
9 Extensions
Our core framework focuses on adaptive, single-variable, perfect interventions in linear non-Gaussian models. Here, we briefly discuss how the proposed method can be extended to more general settings. We defer detailed derivations and proofs to Appendix F.
9.1 Multi-Node Interventions
The proposed method can be generalized to handle simultaneous interventions on a set of variables $E=\{i_{1},\dots,i_{t}\}$. Such an intervention is always informative, as it localizes the row-permutation ambiguity to the intervened set $E$ and effectively disambiguates it from the rest of the graph. Furthermore, if the subgraph induced by the variables in $E$ is acyclic, the internal cycle reversion ambiguity within $E$ is fully resolved. This allows for the unique recovery of the true row for each variable in $E$, enabling the safe parallelization of experiments to accelerate the causal discovery process. We provide a detailed analysis in Appendix F.1.
9.2 Imperfect Interventions
Our approach is also robust to imperfect interventions. The core requirement for identifying a variable’s corresponding row is that the intervention sufficiently perturbs its causal mechanism. As long as the post-intervention row is distinguishable from the set of observational rows after ICA estimation, the intervention is informative. This condition holds for some realistic noisy or incomplete intervention types, enhancing the practical applicability of our method. Further details are discussed in Appendix F.2.
9.3 Generalization to Non-linear Models
While we focused on the linear case, our experimental design strategy could be extended to non-linear structural causal models. Recent advances in non-linear ICA can recover the Jacobian of the model’s inverse function, whose sparsity pattern reveals the underlying causal graph structure up to the same permutation ambiguity found in the linear setting reizinger2023jacobian. Our bipartite matching formulation and reward function could then be applied to resolve this ambiguity. We outline this potential generalization in Appendix F.3.
@@ Appendix A Distribution Equivalence: Formal Definitions, Characterization, and Proofs
To understand how graphs within the same equivalence class differ, we introduce the cycle decomposition of permutations and relate it to structural changes in the graph. While this idea is discussed in [lacerda2012discovering], we include it here for completeness—especially since their treatment is limited to distribution-entailment equivalent LSCMs, rather than distribution-equivalent graphs.
@@ A.4 Example: Matchings in Bipartite Graphs
Here, cross marks indicate non-zero entries in the matrix, and the red entries mark a specific row permutation $\pi$ that selects a unique non-zero element in each row, corresponding to a perfect matching. This selection defines a valid permutation matrix $P_{\pi}$ such that $P_{\pi}(I-W_{\mathrm{ICA}})$ is a valid causal model in the equivalence class with non-zero elements on the diagonal. Figure 3(a) further illustrates how such permutations correspond to bipartite matchings.
@@ B.2 Example: How Interventions Reveal True Matching Edges
As discussed in Proposition 5.1, interventions provide critical information about the true matching in the bipartite representation. Recall that each perfect matching in the bipartite graph $G_{b}=(\{r_{1},\dots,r_{n}\},[n],E)$, where an edge $(r_{i},j)\in E$ exists iff $(I-W_{\mathrm{ICA}})[i,j]\neq 0$, corresponds to a valid row permutation matrix $P_{\pi}$ defining a distribution-equivalent graph.
@@ Appendix C Proofs of Section 6
$f(\mathcal{I},\phi)\;=\;|\Omega|\;-\;\bigl|\Omega^{(\mathcal{I},\phi)}\bigr|$
@@
$\mathcal{I}_{1}=\mathrm{dom}(\psi),\quad\mathcal{I}_{2}=\mathrm{dom}(\psi^{\prime}),\quad\mathcal{I}_{1}\subseteq\mathcal{I}_{2},$
@@
$\Omega_{1}\;=\;\Omega^{(\mathcal{I}_{1},\phi)},\quad\Omega_{2}\;=\;\Omega^{(\mathcal{I}_{2},\phi)},$
@@
and $N_{2}\leq N_{1}$. Let $v\notin\mathcal{I}_{2}$ be the next intervention candidate. In the bipartite‐matching view, assume that $v$ has $k$ possible edges across $\Omega_{1}$. Concretely, recall that each candidate graph in our current equivalence class $\Omega_{1}$ corresponds to a perfect matching in the bipartite graph $G_{b}$ whose left nodes are rows $r_{1},\dots,r_{n}$ and whose right nodes are column positions $1,\dots,n$. Before intervening on variable $v$, the right‐hand node $v$ may be matched to different rows $r_{i}$ across the $\lvert\Omega_{1}\rvert=N_{1}$ graphs. Call these distinct row‐indices $i=z_{1},z_{2},\dots,z_{k}$, so that the set of “candidate edges” for node $v$ is
@@
By the result of Golovin and Krause [golovin2011adaptive], adaptive monotonicity and submodularity guarantee that the adaptive greedy policy achieves at least a (11/e)(1-1/e)-approximation to the optimal policy. ∎